# Invited Paper: CMOS 2.0 - Redefining the Future of Scaling

### v1

@@ Abstract
publicationid: pubid: ©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
As the semiconductor industry advances into the Angstrom era, geometric scaling becomes extremely challenging. Research efforts are increasingly focused on the question: What comes after Complementary FET (CFET) [1]? A relief from the staggering challenges of pure geometric scaling comes from the recent promising advancements in fine pitch 3D bonding [2].Wafer-to-Wafer (W2W) hybrid bonding, as shown in Fig. 1, now enables two paradigm shifts: sub-micron inter-die connectivity and the stacking of heterogeneous Front End Of Line (FEOL) layers with specialized Back End Of Line (BEOL). This opens the path to technology-aware system design exploiting a heterogeneous third dimension.
@@ II-A The CMOS scaling challenge
Finally, the storage of information is distributed across various layers of the compute architecture. From register files and L0 caches inside the processing core up to shared cache resources between processing units, they have different capacity and speed requirements and hence utilize various types of circuitry. A 3D partitioning of the highest cache levels is already implemented in several products today. Going deeper into the lower-level memory structures, critical speed requirements usually limit their capacity. In a fourth partitioning scheme, we propose to disintegrate the memory array from part of its periphery. Encoding/decoding logic can still be implemented in the advanced logic tier, but the bitcells could drive a completely different technology optimization (case D). This partitioning scheme is being studied in some memory technologies and referred to Array-under-CMOS (AuC) [4].
@@ III-B Micro-architectural innovations for 3D
The increased spatial degrees of freedom available in the third dimension can be exploited to increase on-chip memory size, bandwidth, and reduce interconnect latencies not only as PPA boosters in traditional architectures, but even natively targeting architectures to the new implementation fabric. Shared-memory many-PE cluster architectures, such as Mempool and Terapool [8, 9], feature wire-intensive low-diameter interconnects between PEs and many-bank L1 shared memory. These interconnects are extremely unwieldy to implement within a single CMOS layer. Thus, many-PE cluster architectures greatly benefit from the increased routing resources tied to the vertical interconnects available in CMOS 2.0. An initial study was presented in [10] where a Logic-on-Logic (LOL) split was shown to improve interconnect latency, allowing single-cycle access to many memory banks for a large number of PEs, with major reductions in wirelength and power compared to an implementation in standard single-layer CMOS technology.
@@
Compute-in-memory (CIM) architectures are also poised to benefit from CMOS 2.0. CIM has been shown to have great potential for improved efficiency in specific workloads such as dense linear algebra (which dominates deep neural network inference and training). CIM merges computation with storage: it has been extensively explored in research [11], and is getting increased traction in industrial products such as the Metis AIPU [12] from Axelera and EN100 [13] from Encharge AI. CIM’s main challenge is that memory structures enhanced with computing capabilities have significant density overhead compared to standard memories. Intuitively, by exploiting 3D stacking, this overhead can be significantly reduced. For example, [14] proposes a monolithically stacked CFET-based SRAM array, significantly reducing RC parasitics and achieving high energy efficiencies. Additionally, the peripheral circuitry, consisting of data read/write and reduction computations, is implemented in FinFET technology for better latency without impacting the high-density footprint of the SRAM array.
@@ III-C Novel architectural paradigms
Another direction targets the need for supporting multiple data-parallel workloads with high efficiency in hardwired accelerators. A notable example is supporting low arithmetic intensity operators (e.g. matrix-vector multiplication) on systolic arrays primarily designed for matrix-matrix multiplication. This dual-mode operation requires re-configurable interconnects and fine-grained power management of PEs, which is very expensive (in area and power) in traditional CMOS. CMOS 2.0 has the potential to enable architectural flexibility using specialized stacked layers. In the same direction, we could extend the capabilities of the processing elements in the systolic array. Moving from MAC units to more programmable logic requires local storage of control bits (microcode) for cycle-by-cycle operation sequencing. This approach has been tried many times in classical CMOS with coarse-grained reconfigurable arrays (CGRAs) [17]. Unfortunately, CGRAs have had limited commercial success because of their relatively lower efficiency compared to dedicated systolic arrays - in other words, the price paid for flexibility in CGRAs has been too high in standard CMOS.
@@ IV Electronic Design Automation
Products leveraging 2.5D and 3D integration [7] [19] partition the system coarsely into different parts on the subsystem-level, and implement these in different pieces of silicon. The resulting chiplets then communicate through coarse pitch interfaces as indicated in Fig. 1.
@@
Previous works tried to estimate the impact of fine pitch 3D integration on block-level for different schemes like MOL and LOL under various 3D stacking assumptions like Face-to-Face (F2F)/Face-to-Back (F2B) W2W/Die-to-Wafer (D2W) hybrid bonding [20] or monolithic integration [21] [22]. In these, commercial 2D engines are tricked with modified technology files to represent the 3D arrangement of the design. While such approaches can elegantly evaluate some scenarios under certain conditions, they do not represent an expandable foundation for a holistic exploration platform as required for CMOS 2.0.
@@ IV-A Towards EDA for very fine pitch 3D stacks
Assuming a system architecture designed for case 1) and present in the form of its RTL, we believe the subsequent implementation toolchain requires significant extensions. First, when mapping Boolean expressions to gates, the synthesis engine will require the physical context, like the availability of placement and routing resources, as well as the inter-layer parasitics, to improve its QoR. Second, the optimization engines should support cell movements across layers to fully leverage the very fine inter-layer pitch. Third, physical awareness is required to cut the connectivity graph whenever it benefits over pure in-plane routing while satisfying the signal density limits dictated by the respective bonding technologies. With scaling of the bonding pitch, an increasing portion of the system will be able to harness this benefit. These examples highlight the increasing need for tightly coupling synthesis and placement engines to handle the finer interactions along the out-of-plane dimension. Further, the cost functions during placement, like global placement [23] or legalization [24], have to be redesigned. These engines typically leverage estimations of the routing engine; thus, it is crucial to embed characteristics of the complete stack into the parasitics model. Eventually, the router needs adaptation to address critical details such as optimizing routes accessing cell pins now available on both the front-side and back-side of the active layers.
@@ V-A Reliability Wall in Advanced Technology Nodes
A major challenge of continued scaling is test access—the difficulty of accessing hundreds of billions of transistors through only a few hundred to a few thousand pins. This challenge is compounded in 3D heterogeneous integration [25], where multiple wafers are stacked and access points are severely limited [26]. As system complexity grows, minor defects under certain environmental conditions (e.g., supply voltage variations, thermal gradients, process corners) can lead to failures. This has increased reliance on parametric testing and System Level Test (SLT), both of which drive up test complexity and cost [27].
Another critical concern is manufacturing yield. In conventional 3D stacking, the Known Good Die (KGD) approach—bonding only dies that pass wafer probe tests—helps maintain yield [28]. However, in CMOS 2.0 stacking, such pre-selection is not feasible. Without novel solutions, yield can degrade sharply as more layers are bonded, threatening the economic viability of the technology.
Additionally, higher design densities push current and power densities to new extremes, demanding more robust solutions for voltage regulation and heat dissipation [29][30]. Complex interactions—spanning chip placement, system architecture, and hardware–software coordination—further increase the likelihood of physical failures. These factors make it difficult and costly to guarantee defect-free operation in the field, underscoring the need for runtime sensing and monitoring to predict, detect, and mitigate failures dynamically [31].
@@
An increasingly important approach to reliability is Silicon Lifecycle Management (SLM) —a holistic strategy spanning design, manufacturing, deployment, operation, maintenance, and end-of-life [32]. In the design stage, close collaboration between manufacturers and system architects ensures performance, power, and reliability goals are met while accounting for system-level constraints such as power consumption, thermal management, and security. Manufacturing focuses on process control, yield optimization, and quality assurance; deployment ensures seamless system integration; and operation emphasizes continuous monitoring, diagnostics, and maintenance.

### v2

@@ Abstract
publicationid: pubid: ©2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works. DOI: 10.1109/ICCAD66269.2025.11240655
@@ I Introduction
As the semiconductor industry advances into the Angstrom era, geometric scaling becomes extremely challenging. Research efforts are increasingly focused on the question: What comes after Complementary FET (CFET) [cfet2018]? A relief from the staggering challenges of pure geometric scaling comes from the recent promising advancements in fine pitch 3D bonding [chew2024_400nm_pitch].Wafer-to-Wafer (W2W) hybrid bonding, as shown in Fig. 1, now enables two paradigm shifts: sub-micron inter-die connectivity and the stacking of heterogeneous Front End Of Line (FEOL) layers with specialized Back End Of Line (BEOL). This opens the path to technology-aware system design exploiting a heterogeneous third dimension.
@@ II-A The CMOS scaling challenge
Finally, the storage of information is distributed across various layers of the compute architecture. From register files and L0 caches inside the processing core up to shared cache resources between processing units, they have different capacity and speed requirements and hence utilize various types of circuitry. A 3D partitioning of the highest cache levels is already implemented in several products today. Going deeper into the lower-level memory structures, critical speed requirements usually limit their capacity. In a fourth partitioning scheme, we propose to disintegrate the memory array from part of its periphery. Encoding/decoding logic can still be implemented in the advanced logic tier, but the bitcells could drive a completely different technology optimization (case D). This partitioning scheme is being studied in some memory technologies and referred to Array-under-CMOS (AuC) [auc_shairfe].
@@ III-B Micro-architectural innovations for 3D
The increased spatial degrees of freedom available in the third dimension can be exploited to increase on-chip memory size, bandwidth, and reduce interconnect latencies not only as PPA boosters in traditional architectures, but even natively targeting architectures to the new implementation fabric. Shared-memory many-PE cluster architectures, such as Mempool and Terapool [mempool, terapool], feature wire-intensive low-diameter interconnects between PEs and many-bank L1 shared memory. These interconnects are extremely unwieldy to implement within a single CMOS layer. Thus, many-PE cluster architectures greatly benefit from the increased routing resources tied to the vertical interconnects available in CMOS 2.0. An initial study was presented in [mempool3d] where a Logic-on-Logic (LOL) split was shown to improve interconnect latency, allowing single-cycle access to many memory banks for a large number of PEs, with major reductions in wirelength and power compared to an implementation in standard single-layer CMOS technology.
@@
Compute-in-memory (CIM) architectures are also poised to benefit from CMOS 2.0. CIM has been shown to have great potential for improved efficiency in specific workloads such as dense linear algebra (which dominates deep neural network inference and training). CIM merges computation with storage: it has been extensively explored in research [cnm_cim], and is getting increased traction in industrial products such as the Metis AIPU [metisAIPU] from Axelera and EN100 [EN100] from Encharge AI. CIM’s main challenge is that memory structures enhanced with computing capabilities have significant density overhead compared to standard memories. Intuitively, by exploiting 3D stacking, this overhead can be significantly reduced. For example, [3d_cfet_cim] proposes a monolithically stacked CFET-based SRAM array, significantly reducing RC parasitics and achieving high energy efficiencies. Additionally, the peripheral circuitry, consisting of data read/write and reduction computations, is implemented in FinFET technology for better latency without impacting the high-density footprint of the SRAM array.
@@ III-C Novel architectural paradigms
Another direction targets the need for supporting multiple data-parallel workloads with high efficiency in hardwired accelerators. A notable example is supporting low arithmetic intensity operators (e.g. matrix-vector multiplication) on systolic arrays primarily designed for matrix-matrix multiplication. This dual-mode operation requires re-configurable interconnects and fine-grained power management of PEs, which is very expensive (in area and power) in traditional CMOS. CMOS 2.0 has the potential to enable architectural flexibility using specialized stacked layers. In the same direction, we could extend the capabilities of the processing elements in the systolic array. Moving from MAC units to more programmable logic requires local storage of control bits (microcode) for cycle-by-cycle operation sequencing. This approach has been tried many times in classical CMOS with coarse-grained reconfigurable arrays (CGRAs) [CGRA_survey]. Unfortunately, CGRAs have had limited commercial success because of their relatively lower efficiency compared to dedicated systolic arrays - in other words, the price paid for flexibility in CGRAs has been too high in standard CMOS.
@@ IV Electronic Design Automation
Products leveraging 2.5D and 3D integration [amd_3D] [amd_MI300] partition the system coarsely into different parts on the subsystem-level, and implement these in different pieces of silicon. The resulting chiplets then communicate through coarse pitch interfaces as indicated in Fig. 1.
@@
Previous works tried to estimate the impact of fine pitch 3D integration on block-level for different schemes like MOL and LOL under various 3D stacking assumptions like Face-to-Face (F2F)/Face-to-Back (F2B) W2W/Die-to-Wafer (D2W) hybrid bonding [macro_3d_stt] or monolithic integration [cascade2d] [m3dadtco]. In these, commercial 2D engines are tricked with modified technology files to represent the 3D arrangement of the design. While such approaches can elegantly evaluate some scenarios under certain conditions, they do not represent an expandable foundation for a holistic exploration platform as required for CMOS 2.0.
@@ IV-A Towards EDA for very fine pitch 3D stacks
Assuming a system architecture designed for case 1) and present in the form of its RTL, we believe the subsequent implementation toolchain requires significant extensions. First, when mapping Boolean expressions to gates, the synthesis engine will require the physical context, like the availability of placement and routing resources, as well as the inter-layer parasitics, to improve its QoR. Second, the optimization engines should support cell movements across layers to fully leverage the very fine inter-layer pitch. Third, physical awareness is required to cut the connectivity graph whenever it benefits over pure in-plane routing while satisfying the signal density limits dictated by the respective bonding technologies. With scaling of the bonding pitch, an increasing portion of the system will be able to harness this benefit. These examples highlight the increasing need for tightly coupling synthesis and placement engines to handle the finer interactions along the out-of-plane dimension. Further, the cost functions during placement, like global placement [eplace_3d] or legalization [legalizer_3d], have to be redesigned. These engines typically leverage estimations of the routing engine; thus, it is crucial to embed characteristics of the complete stack into the parasitics model. Eventually, the router needs adaptation to address critical details such as optimizing routes accessing cell pins now available on both the front-side and back-side of the active layers.
@@ V-A Reliability Wall in Advanced Technology Nodes
A major challenge of continued scaling is test access—the difficulty of accessing hundreds of billions of transistors through only a few hundred to a few thousand pins. This challenge is compounded in 3D heterogeneous integration [dhananjay2021monolithic], where multiple wafers are stacked and access points are severely limited [marinissen2009testing]. As system complexity grows, minor defects under certain environmental conditions (e.g., supply voltage variations, thermal gradients, process corners) can lead to failures. This has increased reliance on parametric testing and System Level Test (SLT), both of which drive up test complexity and cost [appello2021system].
Another critical concern is manufacturing yield. In conventional 3D stacking, the Known Good Die (KGD) approach—bonding only dies that pass wafer probe tests—helps maintain yield [knickerbocker20122]. However, in CMOS 2.0 stacking, such pre-selection is not feasible. Without novel solutions, yield can degrade sharply as more layers are bonded, threatening the economic viability of the technology.
Additionally, higher design densities push current and power densities to new extremes, demanding more robust solutions for voltage regulation and heat dissipation [hesheng2022_voltage_regulation][oprins2021_cooling_solutions_3d]. Complex interactions—spanning chip placement, system architecture, and hardware–software coordination—further increase the likelihood of physical failures. These factors make it difficult and costly to guarantee defect-free operation in the field, underscoring the need for runtime sensing and monitoring to predict, detect, and mitigate failures dynamically [landman2022applying].
@@
An increasingly important approach to reliability is Silicon Lifecycle Management (SLM) —a holistic strategy spanning design, manufacturing, deployment, operation, maintenance, and end-of-life [tahoori2024silicon]. In the design stage, close collaboration between manufacturers and system architects ensures performance, power, and reliability goals are met while accounting for system-level constraints such as power consumption, thermal management, and security. Manufacturing focuses on process control, yield optimization, and quality assurance; deployment ensures seamless system integration; and operation emphasizes continuous monitoring, diagnostics, and maintenance.